Errata for Selected Papers on Computer Science, first printing (last updated 25 March 1999) (errata to the second printing are not shown here; see WWW)

The running headlines on lefthand pages throughout the book should say Selected Papers on Computer Science (not Selected Papers in Computer Science)! Page 28, line 7 from the bottom Chapter 10 Chapter 11 Page 28, line 3 from the bottom Chapter 13 Chapter 14 Page 54, line 8 from the bottom Deutsche Deutschen Page 55, line 16 Anh. C Anh. C Page 76, line 3 the next page the preceding page Page 81, line 5 the next page page 79 Page 94, line 14 programing programming Page 98, line 13 from the bottom union disjoint union Page 100, bottom line over L over K. of L over K. Page 102, line 12 q · p_{n−1,}. q · p_{n−1,k}. Page 113, line 4 from the bottom Acta Mathematica Acta Informatica Page 115, line 8 from the bottom primarily it primarily because it Page 116, line 4 the former the latter Page 123, line 6 Savillian Savilian Page 123, replacement for lines 10 and 11

[KWEE coan-DIS lai-WAH coan-DIS col-LAYG-i-a DEX-tra, NAY-mo too-AR oo-NAM WEE-kit oo-TRAH-kwe mah-NU.]

Page 127, line 3 Polya Pólya Page 127, replacement for lines 15–18

You who establish colleges with left hand and right, Nobody surpassed one of yours using both hands.

(Thanks to Prof. Mike Wigodsky of Stanford’s Classics Department and Prof. Hendrik Lenstra of Berkeley’s Mathematics Department for hints about translation and pronunciation.) Page 127, bottom line 1974–. 1974–1984. Page 131, line 8 where where we Page 136, line 9 signification significant Page 137, line 23 that there’s a certain that a certain Page 170, line 25 Good Thing, Good Thing. Page 176, line 15 from the bottom they fall they fail Page 176, line 11 from the bottom obso1ete obsolete Page 181, line 7 from the bottom Wright’s Theory Wright, An Introduction to the Theory Page 190, line 8

12 by multiplied 12 be multiplied

Page 202, lines 5 and 6 from the bottom griechis- chen grie- chischen Page 209, line 1 word as word was Page 209, line 17 (abdefg)_{10} (abcdefg)_{10} Page 215, lines 1 and 3 z_{1} z_{l} (twice) Page 227, line 10 from the bottom got hold of a couple got ahold of some Page 259, line 12 1966) . 1966). [Also, a printing glitch has compressed many of the letters on this page.] Page 268, left column Hummurabi Hammurabi (and this entry moves to page 267) Page 269, new entry Lenstra, Hendrik Willem, Jr., 127.

Page 270, left column O’Bierne O’Beirne Page 272, Snell entry Snellius [= Snel van Royen], Willebrord, law, 97.